OpenStudy (anonymous):

In right ∆ABC, m B = 90°, m C = 40°, and BC = 10. What are the other two side lengths of the triangle? AC = 13.1, AB = 8.4 AC = 18.7, AB = 15.8 AC = 14.3, AB = 10.2 AC = 15.6, AB = 11.9

OpenStudy (anonymous):

You will need to first find the third angle, angle a =50 Then use the law of sines to find the other two sides. 50/10=40/x and 50/10= 90/y